OpenStudy (anonymous):

Part 1: Explain how you calculate the restricted value(s) of x for the equation the fraction the binomial x plus 3 over the binomial x plus 5 plus the fraction x over 6 equals the fraction 5 over 6 and what those value(s) represent.

OpenStudy (anonymous):

\[\frac{x+3}{x+5}+\frac{x}{6}=\frac{5}{6}\]Is this where we are starting?

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

The restricted value is x can't be -5; if it is, the first denominator will be zero. Do you need to solve the problem?

OpenStudy (anonymous):

i need to explain which method (cross-multiplication or LCD) you would use to solve the equation and why and also explain the steps to solve the equation using the method you chose in step 2.then i provide the solution to the equation :/ can you help me?

OpenStudy (kinggeorge):

This question should be asked in the mathematics section, and not the feedback section. Thanks.

OpenStudy (anonymous):

oh i didnt notice it was in the feedbck section :(

OpenStudy (anonymous):

I would move the second term to the RHS. That would give\[\frac{x+3}{x+5}=\frac{5-x}{6}\]From there, I would use the cross-multiplication method.

OpenStudy (anonymous):

oh i see thanks :D

OpenStudy (anonymous):

No sweat. Do math every day.